OpenStudy (anonymous):

find the value of c that makes each trinomial a perfect square: x^2-21x+c. I think it's Prime.

OpenStudy (anonymous):

x^2-(42/2)x+((-42)^2) c=(-42)^2

OpenStudy (anonymous):

Please can you show me the steps

OpenStudy (anonymous):

maybe I put it this way x^2-21x+c =(x+a)^2 x^2-21x+c =x^2 +2a +a^2 compare both sides, 2a =-21 a=-21/2 c=a^2 =(-21/2)^2 Sorry that I have made a mistake in my first answer

OpenStudy (anonymous):

Thank you so much